# Concurrent-Allocation Task Execution for Multi-Robot Path-Crossing-Minimal Navigation in Obstacle Environments

### v1

@@ II-C Allocation Matrix
The conditions in (5) and (6) stipulate a basic constraint of the one-to-one correspondence between robots and desired points. There still exists $N!$ candidates to be explored to find the optimal $\bm{\alpha}^{\ast}$, which thus offers flexibility in terms of ordering sequences in MPCM navigation.
@@ II-D Obstacles
The fourth block is a set of $M$ finite-size obstacles $\mathcal{O}_{all}=\{\mathcal{O}_{l}\subseteq\mathbb{R}^{n}~{}|~{}l\in\mathcal{M}\}$ with the index set =$\mathcal{M}=\{1,2,\dots,M\}$.
@@ II-E Minimum-Energy Task Execution
[…]
@@ III Problem formulation
  • P2 (Convergence and maneuvering): All robots converge to their desired points and maneuver with the desired velocity $\mathbb{v}_{d}$ in (3), i.e., $\lim_{t\rightarrow\infty}\{\mathbb{x}_{i}(t)-\mathbb{x}_{k}^{d}(t)\}$ $=0,\lim_{t\rightarrow\infty}\dot{\mathbb{x}}_{i}(t)=\mathbb{v}_{d},i\in\mathbb{Z}_{1}^{N},$ where $k$ is the index for which $\alpha_{i,k}=1$ in (4).
@@
  • A2: The initial distance between arbitrary two desired points satisfies $\|\mathbb{x}_{i}^{d}(0)-\mathbb{x}_{j}^{d}(0)\|\geq r,i\neq j\in\mathcal{V},$ with $r$ given below (2).
  • A3: The initial distance between arbitrary two robots satisfies $\|\mathbb{x}_{i}(0)-\mathbb{x}_{j}(0)\|\geq r,\forall i\neq j\in\mathcal{V}.$
@@
  • A5: The sensing radius $R$ in (2) and safe radius $r$ satisfy $R\geq r+\epsilon,$ with an arbitrary positive constant ϵ+italic-$\epsilon\in\mathbb{R}^{+}$.
@@
A1 rigorously guarantees that $\lim_{t\rightarrow\infty}|\phi(\mathbb{p}_{0}(t))|$ $=0\Rightarrow\lim_{t\rightarrow\infty}\mathrm{dist}(\mathbb{p}_{0}(t),\mathcal{T})=0$. A2 prevents inter-robot collisions if the desired multi-robot navigation is formed. A3 and A4 are necessary conditions for P3 and P4, respectively. A5 is reasonable and intuitive in practice. A6 is a common estimator if a small proportion of robots have access to $\mathbb{v}_{d}$ [41, 31].
@@ IV-A Encoding CBF Constraints
According to Definition 2, we introduce three kinds of tasks, namely, the desired-point convergence, inter-robot collision avoidance, and obstacle avoidance.
@@
where $\delta_{i,k}\in\mathbb{R}^{+}$ denotes the slack variable, ϖ+italic-$\varpi\in\mathbb{R}^{+}$ is the penalty term for $\alpha_{i,k}=0$, $\widehat{\mathbb{v}}_{i}$ represents the $i$-th velocity estimator for $\mathbb{v}_{d}$ in A6, and $\gamma(\cdot)$ is the extended class $\mathcal{K}$ function in Definition 1.
@@
$\frac{\partial\phi_{i,j}^{[2]}(\mathbb{x}_{i})}{\partial\mathbb{x}_{i}^{\mbox{\tiny\sf T}}}(\mathbb{u}_{i}-\widehat{\mathbb{v}}_{i})+\gamma\big{(}\phi_{i,j}^{[2]}(\mathbb{x}_{i})\big{)}\leq 0,\;j\in\mathcal{N}_{i}(t).$ (16)
@@ 20      Execute $\mathbb{u}_{i}^{\ast},\forall i\in\mathcal{V}$ for MPCM navigation;
If the $k$-th desired point is allocated to the $i$-th robot, then the constraints (19e) indicate that $\alpha_{i,k}=1$ and $\alpha_{i,q}=0,\forall q\neq k\in\mathbb{Z}_{1}^{N}$, it follows from (IV-A) that $(1-\alpha_{i,k})\varpi=0$ and $(1-\alpha_{i,q})\varpi=\varpi$, for $q\neq k$, which implies that (19b) can be divided into two subgroups: the $k$-th one is the most strict, while the remaining $N-1$ ones are more relax because of the large penalty $\varpi$, i.e.,
@@ V-A1 Feasibility guarantee
[…]
@@ V-A3 Livelock prevention
where $N\in\mathbb{Z}^{+},m_{i}\in\mathbb{Z}^{+},M\in\mathbb{Z}^{+}$ are the number of robots, neighboring robots (i.e., $m_{i}=|\mathcal{N}_{i}|$) in (2) and obstacles, respectively, and the constraints $\mathbb{g}_{i}^{[1]}\in\mathbb{R}^{N},\mathbb{g}_{i}^{[2]}\in\mathbb{R}^{m_{i}},\mathbb{g}_{i}^{[3]}\in\mathbb{R}^{M},g_{i}^{[4]}\in\mathbb{R}$ are $\mathbb{g}_{i}^{[1]}:=\frac{\partial\Phi_{i}^{[1]}}{\partial\mathbb{x}_{i}^{\mbox{\tiny\sf T}}}(\mathbb{u}_{i}-\widehat{\mathbb{v}}_{i})+\gamma\big{(}\Phi_{i}^{[1]}\big{)}-\mathbb{D}_{i}\varpi-\bm{\delta}_{i},\mathbb{g}_{i}^{[2]}:=\frac{\partial\Phi_{i}^{[2]}}{\partial\mathbb{x}_{i}^{\mbox{\tiny\sf T}}}(\mathbb{u}_{i}-\widehat{\mathbb{v}}_{i})+\gamma\big{(}\Phi_{i}^{[2]}\big{)},\mathbb{g}_{i}^{[3]}:=\frac{\partial\Phi_{i}^{[3]}}{\partial\mathbb{x}_{i}^{\mbox{\tiny\sf T}}}(\mathbb{u}_{i}-\widehat{\mathbb{v}}_{i})+\gamma\big{(}\Phi_{i}^{[3]}\big{)},g_{i}^{[4]}:=\|\mathbb{u}_{i}\|-u_{\max}.$ Here, $\gamma(\cdot)$ is evaluated entrywise, and $\mathbb{D}_{i}:=[1,\cdots,0,\cdots,1]\in\mathbb{R}^{N},\Phi_{i}^{[1]}:=[\phi_{i,1}^{[1]},\cdots,\phi_{i,N}^{[1]}]^{\mbox{\tiny\sf T}}\in\mathbb{R}^{N},\gamma(\Phi_{i}^{[1]}):=[\gamma(\phi_{i,1}^{[1]}),\dots,\gamma(\phi_{i,N}^{[1]})]^{\mbox{\tiny\sf T}}\in\mathbb{R}^{N},\Phi_{i}^{[2]}:=[\phi_{i,j}^{[2]},\cdots,\phi_{i,l}^{[2]}]^{\mbox{\tiny\sf T}}\in\mathbb{R}^{m_{i}},\gamma(\Phi_{i}^{[2]}):=[\gamma(\phi_{i,j}^{[2]}),\dots,$ $,\gamma(\phi_{i,l}^{[2]})]^{\mbox{\tiny\sf T}}\in\mathbb{R}^{m_{i}},\Phi_{i}^{[3]}:=[\phi_{i,1}^{[3]},$$\cdots,\phi_{i,M}^{[3]}]^{\mbox{\tiny\sf T}}\in\mathbb{R}^{M},\gamma(\Phi_{i}^{[3]}):=[\gamma(\phi_{i,1}^{[3]}),\dots,\gamma(\phi_{i,M}^{[3]})]^{\mbox{\tiny\sf T}}\in\mathbb{R}^{M}$ are the column vectors, where $m_{i}=|\mathcal{N}_{i}|\in\mathbb{Z}^{+}$ is the cardinality of the neighbor set $\mathcal{N}_{i}$, and 00 in $\mathbb{D}_{i}$ is at the $k$-th entry such that $\alpha_{i,k}=1$.
@@
$\mathbb{x}_{i}(T^{-1})=\mathbb{x}_{k}^{d}(T^{-1}),\mathbb{u}_{i}(T^{-})=\mathbb{\widehat{v}}_{i}(T^{-1})=\mathbb{v}_{d}(T^{-1}),$ (28)
@@
$\delta_{i,k}(T^{-1})=0,\delta_{i,q}(T^{-1})=0,\forall q\neq k\in\mathbb{Z}_{1}^{N},$ (29)
@@ V-B Decoupling and Analysis of P1-P4
Combining (V-B) and (33), one has that the CATE algorithm (19) will degenerate into the minimal-energy task execution (27) if $t>T_{1}$, which can guide the robots towards the allocated desired points.
The livelock scenario where the optimal $\bm{\alpha}^{\ast}$ is changed but still satisfies $\|\mathbb{1}_{N}^{\mbox{\tiny\sf T}}\bm{\alpha}^{\ast}(t)-\mathbb{1}_{N}^{\mbox{\tiny\sf T}}\|=0,\forall t>T_{1}$, is well prevented by Lemma 2. Therefore, there will eventually be only one optimal $\bm{\alpha}^{\ast}$ satisfying $\|\mathbb{1}_{N}^{\mbox{\tiny\sf T}}\bm{\alpha}^{\ast}(t)-\mathbb{1}_{N}^{\mbox{\tiny\sf T}}\|=0$ for $t\geq T_{1}$. The proof is thus completed. ∎
@@ V-B1 Proof of P3, P4
If the input constraints are not activated (i.e., $g_{i}^{[4]}<0$), then $g_{i}^{[4]}\leq 0$ in (27) can be omitted. Therefore, the optimization (27) becomes a typical control-lyapunov-function-control-barrier-function quadratic program (CLF-CBF QP) problem [37], namely,
@@
It follows from Eqs. (IV-A) and (IV-A) that ${\partial\phi_{i,j}^{[2]}(\mathbb{x}_{i})}/{\partial\mathbb{x}_{i}}={-(\mathbb{x}_{i}-\mathbb{x}_{j})}/{\|\mathbb{x}_{i}-\mathbb{x}_{j}\|}\neq\mathbf{0}_{2},\forall\mathbb{x}_{i}\in\mathcal{T}_{i,j}^{[2]}$ and ${\partial\phi_{i,l}^{[3]}(\mathbb{x}_{i})}/{\partial\mathbb{x}_{i}}$ $={-(\mathbb{x}_{i}-\mathbb{x}_{l}^{o})}/{\|\mathbb{x}_{i}-\mathbb{x}_{l}^{o}\|}\neq\mathbf{0}_{2},\forall x_{i}\in\mathcal{T}_{i,l}^{[3]}$ are both locally Lipschitz continuous. One has that $\mathbb{u}_{i}^{\ast}$ in (34) are locally Lipschitz continuous according to (Theorem 3 in [37]). The proof is almost the same as [37] which is thus omitted here. Combining two cases together yields $\mathbb{u}_{i}^{\ast}$ for (27) are locally Lipschitz continuous. The proof is thus completed. ∎
@@
Figure 6: First simulation: Column formation in obstacle environments. (a) Trajectories of ten robots forming a column formation in the presence of two obstacles, using fixed-ordering task execution (FOTE) [30]. (b) Trajectories of ten robots from the same initial setting to a column formation, using DAPT [13] + ORCA). (c) Trajectories of ten robots from the same initial setting to a column formation, using CAPT [15] + ORCA. (d) Trajectories of ten robots from the same initial setting to a column formation, using the proposed CATE. The blue, and yellow triangles and the dashed lines denote the initial, final positions and trajectories of the robots, respectively. The orange irregular polygons, red circles, and dashed ellipses are the obstacles, disk boundaries, and region of path crossings, respectively. Note that the green labels in subfigures (c) represent the switching of the desired-point assignments compared with the original red labels in subfigure (b). The blue labels in subfigure (d) denote the differences of the desired-point assignments compared with subfigure (c).
@@ V-B2 Proof of P2
Under A2 and A6, all robots governed by (1) and (27) can finally converge to the desired points and maneuver with $\mathbb{v}_{d}$ in (3), i.e., P2 is achieved.
@@ V-B3 Proof of P1
Since the path crossings $\mathcal{C}$ in P1 are generally defined to be the same positions of possibly different robots at possibly different historical time instances, which can only be calculated after a long period of time or once the formation maneuvering has been completed. Consequently, it is difficult to calculate and minimize the path crossings in real time for the robot group $\mathcal{V}$, which thus motivates the CATE algorithm (19) to indirectly and implicitly minimize possible path crossings by minimizing the desired-point allocation cost $b\|\mathbb{1}_{N}^{\mbox{\tiny\sf T}}\bm{\alpha}-\mathbb{1}_{N}^{\mbox{\tiny\sf T}}\|^{2}$ and the slack variables $\|\bm{\delta}_{i}\|^{2}$ simultaneously. Precisely, first, the allocation cost $\|\mathbb{1}_{N}^{\mbox{\tiny\sf T}}\bm{\alpha}-\mathbb{1}_{N}\|=0$ and the integer constraints (19d) ensure $N!$ choices of $\bm{\alpha}$ for an arbitrary one-to-one correspondence between desired points and robots. Then, the on-the-fly minimization of $\|\bm{\delta}_{i}\|^{2}$ further selects the most appropriate one-to-one correspondence among $N!$ choices between robots and their desired points when the obstacles occupy the passing space. In this way, the possible path crossings can be implicitly minimized by the flexible-ordering coordination in obstacle environments, which increases the efficiency of achieving MPCM navigation.
@@
According to the definition of hardness [49], the hardness characterization is commonly applicable for decision-making problems, which only involve integer variables, such as the shortest path problem with integer linear programming (ILP) [50]. However, such a definition is not suitable for the MPCM navigation in Problem 1, because Problem 1 not only considers the robot-point-allocation but also involves the robot motion control, which inevitably contains integer and real variables $\bm{\alpha},\mathbb{u}_{i}$. As a remedy, we analyze Problem 1 in a different way to roughly represent its hardness. Specifically, we first consider the basic one-to-one matching problem between robots and the desired points, which is a bipartite matching problem [49]. It has been shown in [13] that such a matching problem can be solved in polynomial time by the Hungarian algorithm. However, since Problem 1 contains additional complex constraints of robot maneuver and convergence, collision, and obstacle avoidance, it significantly intensifies the hardness of Problem 1. Moreover, to better understand the computational complexity of Problem 1 using the proposed CATE algorithm (19), please refer to Remark 7.
Remark 7.
According to [51], the computational complexities of the CATE algorithm (19) for each robot $i$ are calculated to be $O_{i}(2^{N^{2}})$, where $N^{2}$ represents the number of integer variables (i.e., the allocation matrix $\bm{\alpha}\in\mathbb{R}^{N\times N}$). Note that the computational complexities $O_{i}$ grow exponentially with the increasing number of robots $N$ (i.e., the integer variables $N^{2}$ will increase as well), making the problem challenging to solve and extend to large-scale scenarios. Additionally, such limitations and future work on a fully distributed algorithm to reduce computation burden will be summarized in Section VII-B later. It is still worth mentioning that once the integer variables are removed, the CATE optimization (19) will degenerate to the minimal-energy task execution problem (27), which becomes a control-lyapunov-function-control-barrier-function quadratic program (CLF-CBF QP) problem and is generally convex [37].
Remark 8.
For the common undesirable deadlock problem where the trajectory of the robot is directed at some point toward the center of the obstacles [52], the proposed CATE algorithm (19) can conveniently deal with it by reallocating the desired points to different robots to change the direction of the robot’s trajectory. Precisely, suppose the robot $i$, allocated to the desired point $k$, gets stuck and stops in a deadlock at some point due to the obstacles. It follows from Lemma 5 of the revised version the remaining robots will keep moving toward their allocated desired points by reducing the slack vectors $\|\bm{\delta}_{i}\|^{2}$ in the cost functions (19). However, after some time, since robot $i$ remains stuck and the slack variable $\delta_{i,k}$ remains unchanged, the cost function in (19) can no longer decrease. In such a situation, since the proposed CATE algorithm (19) is updated on the fly, it will iteratively find a new allocation matrix $\bm{\alpha}$ to assign a different desired point $l$ to robot $i$ to further decrease the cost functions in (19). Therefore, the original direction of the robot $i$ will be changed, and it is no longer toward the center of obstacles, and thus robot $i$ breaks the original symmetry and escapes the undesirable deadlock situations, which implies that robot $i$ can continue the navigation task without interruption.
VI Simulations and Experiments
@@
TABLE I: Second simulation: Statistical comparison of the “Arrow formation” using different methods in 16 different robot-obstacle group settings.

### v2

@@ II-C Allocation Matrix
The conditions in (5) and (6) stipulate a basic constraint of the one-to-one correspondence between robots and desired points. There still exist $N!$ candidates to be explored to find the optimal $\bm{\alpha}^{\ast}$, which thus offers flexibility in terms of ordering sequences in MPCM navigation.
@@ II-D Obstacles
The fourth block is a set of $M$ finite-size obstacles $\mathcal{O}_{all}=\{\mathcal{O}_{l}\subseteq\mathbb{R}^{n}~|~l\in\mathcal{M}\}$ with the index set $\mathcal{M}=\{1,2,\dots,M\}$.
@@ II-E Minimum-Energy Task Execution
where $h(\mathbb{x}_{i})=-\phi(\mathbb{x}_{i})$ satisfying the constraint in (9b) is a CBF [37], $\bm{\delta}_{i}\in\mathbb{R}^{+}$ is a slack variable to measure the extent of task violation, and $\gamma(\cdot):(-b,a)\rightarrow\mathbb{R}$ is an extended class $\mathcal{K}$ function with $a,b\in\mathbb{R}^{+}$, which is strictly increasing and $\gamma(0)=0$ [37]. Here, $\gamma\big(h(\mathbb{x}_{i})\big)$ can regulate how fast the robots approach the boundary of the target set in (8).
@@ III Problem formulation
  • P2 (Convergence and maneuvering): All robots converge to their desired points and maneuver with the desired velocity $\mathbb{v}_{d}$ in (3), i.e., $\lim_{t\rightarrow\infty}\{\mathbb{x}_{i}(t)-\mathbb{x}_{k}^{d}(t)\}$ $0,\lim_{t\rightarrow\infty}\dot{\mathbb{x}}_{i}(t)=\mathbb{v}_{d},i\in\mathbb{Z}_{1}^{N},$ where $k$ is the index for which $\alpha_{i,k}=1$ in (4).
@@
  • A2: The initial distance between arbitrary two desired points satisfies $\|\mathbb{x}_{i}^{d}(0)-\mathbb{x}_{j}^{d}(0)\|>r,i\neq j\in\mathcal{V},$ with $r$ given below (2).
  • A3: The initial distance between arbitrary two robots satisfies $\|\mathbb{x}_{i}(0)-\mathbb{x}_{j}(0)\|>r,\forall i\neq j\in\mathcal{V}.$
@@
  • A5: The sensing radius $R$ in (2) and safe radius $r$ satisfy $R\geq r+\epsilon,$ with an arbitrary positive constant $\epsilon\in\mathbb{R}^{+}$.
@@
A1 rigorously guarantees that $\lim_{t\rightarrow\infty}|\phi(\mathbb{p}_{0}(t))|$ $0\Rightarrow\lim_{t\rightarrow\infty}\mathrm{dist}(\mathbb{p}_{0}(t),\mathcal{T})=0$. A2 prevents inter-robot collisions if the desired multi-robot navigation is formed. A3 and A4 are necessary conditions for P3 and P4, respectively. A5 is reasonable and intuitive in practice. A6 is a common estimator if a small proportion of robots have access to $\mathbb{v}_{d}$ [41, 31].
@@ IV-A Encoding CBF Constraints
where $\delta_{i,k}\in\mathbb{R}^{+}$ denotes the slack variable, $\varpi\in\mathbb{R}^{+}$ is the penalty term for $\alpha_{i,k}=0$, $\widehat{\mathbb{v}}_{i}$ represents the $i$-th velocity estimator for $\mathbb{v}_{d}$ in A6, and $\gamma(\cdot)$ is the extended class $\mathcal{K}$ function in Definition 1.
@@
$\frac{\partial\phi_{i,j}^{[2]}(\mathbb{x}_{i})}{\partial\mathbb{x}_{i}^{\mbox{\tiny\sf T}}}(\mathbb{u}_{i}-\widehat{\mathbb{v}}_{i})+\gamma\big(\phi_{i,j}^{[2]}(\mathbb{x}_{i})\big)\leq 0,\;j\in\mathcal{N}_{i}(t).\$(16)
@@ IV-B Constrained Optimization
If the $k$-th desired point is allocated to the $i$-th robot, then the constraints (19e) indicate that $\alpha_{i,k}=1$ and $\alpha_{i,q}=0,\forall q\neq k\in\mathbb{Z}_{1}^{N}$, it follows from (IV-A) that $(1-\alpha_{i,k})\varpi=0$ and $(1-\alpha_{i,q})\varpi=\varpi$, for $q\neq k$, which implies that (19b) can be divided into two subgroups: the $k$-th one is the most strict, while the remaining N1N-1 ones are more relax because of the large penalty $\varpi$, i.e.,
@@ V-A1 Feasibility guarantee
$\big\{$ $\mathbb{u}_{i}=\widehat{\mathbb{v}}_{i},\delta_{i,k}\mbox{ is sufficiently large},$
@@ V-A3 Livelock prevention
where $N\in\mathbb{Z}^{+},m_{i}\in\mathbb{Z}^{+},M\in\mathbb{Z}^{+}$ are the number of robots, neighboring robots (i.e., $m_{i}=|\mathcal{N}_{i}|$) in (2) and obstacles, respectively, and the constraints $\mathbb{g}_{i}^{[1]}\in\mathbb{R}^{N},\mathbb{g}_{i}^{[2]}\in\mathbb{R}^{m_{i}},\mathbb{g}_{i}^{[3]}\in\mathbb{R}^{M},g_{i}^{[4]}\in\mathbb{R}$ are $\mathbb{g}_{i}^{[1]}:=\frac{\partial\Phi_{i}^{[1]}}{\partial\mathbb{x}_{i}^{\mbox{\tiny\sf T}}}(\mathbb{u}_{i}-\widehat{\mathbb{v}}_{i})+\gamma\big(\Phi_{i}^{[1]}\big)-\mathbb{D}_{i}\varpi-\bm{\delta}_{i},\mathbb{g}_{i}^{[2]}:=\frac{\partial\Phi_{i}^{[2]}}{\partial\mathbb{x}_{i}^{\mbox{\tiny\sf T}}}(\mathbb{u}_{i}-\widehat{\mathbb{v}}_{i})+\gamma\big(\Phi_{i}^{[2]}\big),\mathbb{g}_{i}^{[3]}:=\frac{\partial\Phi_{i}^{[3]}}{\partial\mathbb{x}_{i}^{\mbox{\tiny\sf T}}}(\mathbb{u}_{i}-\widehat{\mathbb{v}}_{i})+\gamma\big(\Phi_{i}^{[3]}\big),g_{i}^{[4]}:=\|\mathbb{u}_{i}\|-u_{\max}.$ Here, $\gamma(\cdot)$ is evaluated entrywise, and $\mathbb{D}_{i}:=[1,\cdots,0,\cdots,1]\in\mathbb{R}^{N},\Phi_{i}^{[1]}:=[\phi_{i,1}^{[1]},\cdots,\phi_{i,N}^{[1]}]^{\mbox{\tiny\sf T}}\in\mathbb{R}^{N},\gamma(\Phi_{i}^{[1]}):=[\gamma(\phi_{i,1}^{[1]}),\dots,\gamma(\phi_{i,N}^{[1]})]^{\mbox{\tiny\sf T}}\in\mathbb{R}^{N},\Phi_{i}^{[2]}:=[\phi_{i,j}^{[2]},\cdots,\phi_{i,l}^{[2]}]^{\mbox{\tiny\sf T}}\in\mathbb{R}^{m_{i}},\gamma(\Phi_{i}^{[2]}):=[\gamma(\phi_{i,j}^{[2]}),\dots,$ $\gamma(\phi_{i,l}^{[2]})]^{\mbox{\tiny\sf T}}\in\mathbb{R}^{m_{i}},\Phi_{i}^{[3]}:=[\phi_{i,1}^{[3]},$ $\cdots,\phi_{i,M}^{[3]}]^{\mbox{\tiny\sf T}}\in\mathbb{R}^{M},\gamma(\Phi_{i}^{[3]}):=[\gamma(\phi_{i,1}^{[3]}),\dots,\gamma(\phi_{i,M}^{[3]})]^{\mbox{\tiny\sf T}}\in\mathbb{R}^{M}$ are the column vectors, where $m_{i}=|\mathcal{N}_{i}|\in\mathbb{Z}^{+}$ is the cardinality of the neighbor set $\mathcal{N}_{i}$, and 0 in $\mathbb{D}_{i}$ is at the $k$-th entry such that $\alpha_{i,k}=1$.
@@
$\mathbb{x}_{i}(T^{-})=\mathbb{x}_{k}^{d}(T^{-}),\mathbb{u}_{i}(T^{-})=\mathbb{\widehat{v}}_{i}(T^{-})=\mathbb{v}_{d}(T^{-}),$ (28)
@@
$\delta_{i,k}(T^{-})=0,\delta_{i,q}(T^{-})=0,\forall q\neq k\in\mathbb{Z}_{1}^{N},$ (29)
@@ V-B Decoupling and Analysis of P1-P4
Combining (V-B) and (33), one has that the CATE algorithm (19) will degenerate into the minimal-energy task execution (27) if $t>T_{1}$, which can guide the robots towards the allocated desired points. The livelock scenario where the optimal $\bm{\alpha}^{\ast}$ is changed but still satisfies $\|\mathbb{1}_{N}^{\mbox{\tiny\sf T}}\bm{\alpha}^{\ast}(t)-\mathbb{1}_{N}^{\mbox{\tiny\sf T}}\|=0,\forall t>T_{1}$, is well prevented by Lemma 2. Therefore, there will eventually be only one optimal $\bm{\alpha}^{\ast}$ satisfying $\|\mathbb{1}_{N}^{\mbox{\tiny\sf T}}\bm{\alpha}^{\ast}(t)-\mathbb{1}_{N}^{\mbox{\tiny\sf T}}\|=0$ for $t\geq T_{1}$. The proof is thus completed. ∎
@@ V-B1 Proof of P3, P4
If the input constraints are not activated (i.e., $g_{i}^{[4]}<0$), then $g_{i}^{[4]}\leq 0$ in (27) can be omitted. The optimization (27) becomes a control-lyapunov-function-control-barrier-function quadratic program (CLF-CBF QP) problem [37], namely,
@@
It follows from Eqs. (IV-A) and (IV-A) that ${\partial\phi_{i,j}^{[2]}(\mathbb{x}_{i})}/{\partial\mathbb{x}_{i}}={-(\mathbb{x}_{i}-\mathbb{x}_{j})}/{\|\mathbb{x}_{i}-\mathbb{x}_{j}\|}\neq\mathbf{0}_{2},\forall\mathbb{x}_{i}\in\mathcal{T}_{i,j}^{[2]}$ and ${\partial\phi_{i,l}^{[3]}(\mathbb{x}_{i})}/{\partial\mathbb{x}_{i}}$ ${-(\mathbb{x}_{i}-\mathbb{x}_{l}^{o})}/{\|\mathbb{x}_{i}-\mathbb{x}_{l}^{o}\|}\neq\mathbf{0}_{2},\forall x_{i}\in\mathcal{T}_{i,l}^{[3]}$ are both locally Lipschitz continuous. One has that $\mathbb{u}_{i}^{\ast}$ in (34) are locally Lipschitz continuous according to (Theorem 3 in [37]). The proof is almost the same as [37] which is thus omitted here. Combining two cases together yields $\mathbb{u}_{i}^{\ast}$ for (27) are locally Lipschitz continuous. The proof is thus completed. ∎
@@ V-B2 Proof of P2
Under A2 and A6, all the robots governed by (1) and (27) can finally converge to the desired points and maneuver with $\mathbb{v}_{d}$ in (3), i.e., P2 is achieved.
@@ V-B3 Proof of P1
Since the path crossings $\mathcal{C}$ in P1 are generally defined to be the same positions of possibly different robots at possibly different historical time instances, which can only be calculated after a long period of time or once the formation maneuvering has been completed. Consequently, it is difficult to calculate and minimize the path crossings in real time for the robot group $\mathcal{V}$, which thus motivates the CATE algorithm (19) to indirectly and implicitly minimize possible path crossings by minimizing the desired-point allocation cost $b\|\mathbb{1}_{N}^{\mbox{\tiny\sf T}}\bm{\alpha}-\mathbb{1}_{N}^{\mbox{\tiny\sf T}}\|^{2}$ and the slack variables $\|\bm{\delta}_{i}\|^{2}$ simultaneously. Precisely, first, the allocation cost $\|\mathbb{1}_{N}^{\mbox{\tiny\sf T}}\bm{\alpha}-\mathbb{1}_{N}\|=0$ and the integer constraints (19d) ensure $N!$ choices of $\bm{\alpha}$ for an arbitrary one-to-one correspondence between desired points and robots. Then, the on-the-fly minimization of $\|\bm{\delta}_{i}\|^{2}$ further selects the most appropriate one-to-one correspondence among $N!$ choices between robots and their desired points when the obstacles occupy the passing space. In this way, the possible path crossings can be implicitly minimized.
@@
For the common undesirable deadlock problem where the trajectory of the robot is directed at some point toward the center of the obstacles [49], the proposed CATE algorithm (19) can conveniently deal with it by reallocating the desired points to different robots to change the direction of the robot’s trajectory. Precisely, suppose the robot $i$, allocated to the desired point $k$, gets stuck and stops in a deadlock at some point due to the obstacles. It follows from Lemma 5 that the remaining robots will keep moving toward their allocated desired points by reducing the slack vectors $\|\bm{\delta}_{i}\|^{2}$ in the cost functions (19). However, after some time, since robot $i$ remains stuck and the slack variable $\delta_{i,k}$ remains unchanged, the cost function in (19) can no longer decrease. In such a situation, since the proposed CATE algorithm (19) is updated on the fly, it will iteratively find a new allocation matrix $\bm{\alpha}$ to assign a different desired point $l$ to robot $i$ to further decrease the cost functions in (19). Therefore, the original direction of the robot $i$ will be changed, and it is no longer toward the center of obstacles, and thus robot $i$ breaks the original symmetry and escapes the undesirable deadlock situations, which implies that robot $i$ can continue the navigation task without interruption.
@@
TABLE I: Summary of the problem’s hardness in unlabeled multi-robot motion planning works.
Problem categrary Hardness Robot dynamics Obstacles Optimization objective Dimension Positioning of robot, deseired points and obstacles
Unlabelled GMP[50, 51] PSPACE-hard [52] Static 2D No requirement
Unlabelled MPL-F[53, 54, 55] P Shortest path length 2D Well separation
Unlabelled MPL-O[56] P Static Path-length minimization 2D Well separation
Problem 1 Undetermined Static/moving Path-crossing minimization 2D/3D Assumptions A2-A4
GMP: General motion planning, MPL-F: Minimum-path-length motion planning in free space, MPL-O: Minimum-path-length motion planning in obstacle environments.
Figure 6: First simulation: Column formation in obstacle environments. (a) Trajectories of ten robots forming a column formation in the presence of two obstacles, using fixed-ordering task execution (FOTE) [30]. (b) Trajectories of ten robots from the same initial setting to a column formation, using DAPT [13] + ORCA. (c) Trajectories of ten robots from the same initial setting to a column formation, using CAPT [15] + ORCA. (d) Trajectories of ten robots from the same initial setting to a column formation, using the proposed CATE. The blue and yellow triangles and the dashed lines denote the initial, final positions, and trajectories of the robots, respectively. The orange irregular polygons, red circles, and dashed ellipses are the obstacles, disk boundaries, and region of path crossings, respectively. Note that the green labels in subfigure (c) represent the switching of the desired-point assignments compared with the original red labels in subfigure (b). The blue labels in subfigure (d) denote the differences of the desired-point assignments compared with subfigure (c).
V-C Complexity Discussion of Problem 1
Due to the obstacles occupying the passing space between robots and their desired points, and the requirements for adapting to dynamic environments, it is not feasible to assign fixed costs for minimizing path-crossing points for each robot like the classic 1-1 matching problem [15, 16]. To address this, we introduce the allocation cost $\|\mathbb{1}_{N}^{\mbox{\tiny\sf T}}\bm{\alpha}-\mathbb{1}_{N}^{\mbox{\tiny\sf T}}\|$ together with the constraint in (19e) that simply restricts the solution space to the $N!$ possible one-to-one mappings, rather than computing the optimal assignment directly. Subsequently, the additional costs of slack variable $\|\bm{\delta}_{i}\|^{2}$ further select the most appropriate assignment. Although the cost $\|\mathbb{1}_{N}^{\mbox{\tiny\sf T}}\bm{\alpha}-\mathbb{1}_{N}^{\mbox{\tiny\sf T}}\|$ is influenced by the allocation of other robots, the common objective of the one-to-one correspondence in (5) and (6) remains unaffected. Moreover, this cost also offers flexibility and adaptability in matching, particularly in obstacles and dynamic environments. However, in contrast to [15, 16], this cost design makes it difficult to determine whether the matching problem is efficiently solvable. Currently, the CATE algorithm (19) is formulated to be the MIQP framework that is solved using the Gurobi optimization tool [43]. Furthermore, as shown by the shortest convergence time of our algorithm in Table II, the matching problem is solved with favorable performance. The theoretical analysis of whether the matching in our problem is optimal or not will be a future direction.
[b]
TABLE II: Second simulation: Statistical comparison of the “Arrow formation” using different methods in 16 different robot-obstacle group settings.